X mx^mx^mx....=y^ny^ny^ny ..., then dy/dx is equal to

Xmx^mx^mx....=y^ny^ny^ny
..., then dy/dx is equal to

Grade:12

1 Answers

Mohammed Kasib Siddiqui
5 years ago
Well the explanation will be a little tricky. I dont know that you will under stand it or not 
On both the sides the power os going till infinite.
So we can reduce the expression as-
x^my^ny^ny^ny........= y^ny^ny^ny......
(x^m/y^n)^ny^ny^ny....=1
Now we can easily remove the power. 
x^m/y^n=1
x^m=y^n
Differentiate..
mx^m-1=ny^n-1 dy/dx
reaarange to get dy/dx
 
plz approve i have worked alot on this ques.